OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5?

OpenStudy (anonymous):

A. (1, 7) B. (5, 0) C. (2, -1) D. (0, 5)

OpenStudy (anonymous):

Is this where I plug the numbers in and try to get them to equal the same, or is this different?

OpenStudy (anonymous):

.substitute and see which works :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yup. plugging is one true and tried method :P

OpenStudy (mertsj):

Did you get D?

OpenStudy (anonymous):

Okay so I don't get this.. what am I doing wrong? The equations do not equal up. Here is what I got: y = -3(1)+5 = 6 y = 2(7)+5 = 12 y = -3(5)+5 = 10 y = 2(0)+5 = 5 y = -3(2)+5 = 7 y = 2(-1)+5 = 4 y = -3(0)+5 = 5 y = 2(5)+5 = 10

OpenStudy (mertsj):

y = -3(0)+5 = 5 y = 2(5)+5 = 10 The trouble is that x = 0 and you replaced x with 5 in the second part.

OpenStudy (anonymous):

Oh! Alright. So my answer would be D then?

OpenStudy (mertsj):

That's what I got.

OpenStudy (anonymous):

It makes sense to me too. Thanks Mertsj! :)

OpenStudy (mertsj):

yw